mass of granulated sugar ( C12H22O11) in grams. 14.87 grams



mass in kg
mass in cg
moles of sugar
molecules of sugar
atoms of carbon present
atoms of hydrogen present
atoms of oxygen present
Answer #1

mass of sample of sugar = 14.87 g

1)

1 kg = 1000g or 1g = 10-3 kg

thus mass= 14.87x 10-3 kg  = 1.487x10-2 kg

2) cg is centigram

1cg = 10-2 cg

thus mass of sample = 14.87x10-2cg = 0.1487 cg

3) moles = mass/ molar mass

molar mass of sugar = 342g/mol

Thus moles of sugar in sample = 14.87g /342g/mol = 0.04347 mol

4)1 molecule of sugar has 12 atoms from formula C12 H22 O11

molecules of sugar in sample = moles x Avogardo sumber

= 0.04347 mol x 6.022 x1023

Atoms of C = 12 x  0.04347 mol x 6.022 x1023 = 3.14x1021

5) atoms of H present = molecules x 22 = 0.04347 mol x 6.022 x1023 x22=5.75x1023

6) atoms of O = molecules x 11 = 0.04347 mol x 6.022 x1023 x11 =2.88x1023